How many grams of diethyl ether are formed from 0.55 mole of ethanol?


1
Expert's answer
2021-02-01T03:59:21-0500

2C2H5OH → (C2H5)2O + H2O

According to the reaction:

n((C2H5)2O) = "\frac{1}{2}" n(C2H5OH) = "\frac{1}{2} \times 0.55 = 0.275 \;mol"

M((C2H5)2O) = 74.12 g/mol

"m = n \times M"

m((C2H5)2O) "= 0.275 \times 74.12 = 20.38 \;g"

Answer: 20.38 g